OpenStudy (anonymous):

What is an algebraic expression that supports a number subtracted from 8 is equal to 4

OpenStudy (aum):

Let the number be x. 8 - x = 4

OpenStudy (anonymous):

OK thank you so much!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

I have another one, I am not good at these. Which of the following algebraic equations models the English sentence, a number decreased by two is equal to four?

OpenStudy (anonymous):

@aum @blahhedup101

OpenStudy (anonymous):

is this one of those multiple choice ?

OpenStudy (anonymous):

No, just a random equation that will fit the description.

OpenStudy (anonymous):

oh ... well i found this online w − 2 = 4

OpenStudy (anonymous):

Can you help me on one more? I will become your fan if you do.

OpenStudy (anonymous):

umm sure

OpenStudy (anonymous):

What sentence could best describe 4/p=2?

OpenStudy (aum):

Four divided by a number is equal to 2.

OpenStudy (anonymous):

Thanks! it was right. I will fan both of you.

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

np.. are you taking algebra 1 online?

OpenStudy (anonymous):

Yes, why?

OpenStudy (anonymous):

hold on imma message you

OpenStudy (anonymous):

ok